OpenStudy (anonymous):

PLEASE HELP ME I WILL FAN AND MEDAL!!! A line passes through the origin and the point (2,6). The x-axis represents time in minutes, and the y-axis represents gallons. What rate does the slope represent? A. 12 minutes per gallon B. 3 gallons per minute C. 2 gallons per 6 minutes D. 3 minutes per gallon

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

origin is (0,0).....x1 = 0 and y1 = 0 (2,6)...x2 = 2 and y2 = 6 slope(m) = (y2 - y1) / (x2 - x1) now we sub slope(m) = (6 - 0) / (2 - 0) slope(m) = 6/2 which reduces to 3/1 3 gallons per minute

OpenStudy (anonymous):

Thank you. 2 more questions? please?

OpenStudy (texaschic101):

I can try...what ya got

OpenStudy (texaschic101):

just tag me